1. Find the equation of the line passing through the point (3,4) that is perpendicular to the line passing through the points (7,9) and (8,1). Put the answer in point slope form.
2. Put the equation from problem 1 in slope-intercept form
3. Put the equation from problem 1 in standard form
4. Find the x and y- intercepts of the line you found in problem 1

Answer by stanbon(75887)

1. Find the equation of the line passing through the point (3,4) that is perpendicular to the line passing through the points (7,9) and (8,1). Put the answer in point slope form.
m=[1-9]/[8-7]=-8
So the line perpendicular has slope 1/8 and and passes thru (3,4)
4=(1/8)3+b
b=4-(3/8)
b=29/8
y=(1/8)x+(29/8)
-----------------------
2. Put the equation from problem 1 in slope-intercept form
It is.
----------------
3. Put the equation from problem 1 in standard form
x-8y=-29
----------------------
4. Find the x and y- intercepts of the line you found in problem 1
Let x=0, then y=29/8 ; that is the y-intercept
Let y=0, then x=-29
---------------
Cheers,
Stan H.